OpenStudy (anonymous):

A new track in an amusement park will be elliptical with center (0,0), equation x^2/2500+y^2/8100=1 models path of the track. Units in yards. how long is the major axis of the track. explain how you found the distance.

OpenStudy (anonymous):

Remember that our form for a vertical ellipse is \(\frac{(x-h)^2}{b^2}+\frac{(y-k)^2}{a^2}=1\) where \(a>b\) and \(a\), \(b\) correspond to the lengths of our major (horizontal) and minor (vertical) axes, respectively. In our case, we have \(a=90\), \(b=50\); this means our major axis is 90 yd long.

OpenStudy (anonymous):

I thought that a is 90 and then2a=180 yds long.